# One Shot Dominance: Knowledge Poisoning Attack on Retrieval-Augmented Generation Systems

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and found widespread applications in daily life. However, despite their impressive abilities, LLMs still face challenges such as outdated knowledge and hallucination issues as knowledge continues to evolve Achiam et al. (2023); Touvron et al. (2023); Anil et al. (2023). To address these limitations, researchers have introduced the Retrieval-Augmented Generation (RAG) approach to improve LLMs Tu et al. (2024); Zhao et al. (2024). This technology has been widely adopted in both industrial applications and academic research Al Ghadban et al. (2023); Wang et al. (2024a); Loukas et al. (2023); Kumar et al. (2023); Prince et al. (2024).
@@
Although RAG significantly improves the performance of the LLM response, it introduces potential security vulnerabilities. The security of RAG systems is influenced by both the inherent safety mechanisms of LLMs Tan et al. (2024); Xue et al. (2024); Chaudhari et al. (2024) and the potential manipulation of external context Zou et al. (2024); Zhang et al. (2024). Attackers can attempt to mislead LLMs into generating incorrect outputs by poisoning knowledge bases with malicious content. This attack surface is particularly concerning as knowledge bases are often the most accessible component of RAG systems. For example, when RAG systems utilize public resources like Wikipedia Thakur et al. (2021) for current events information Chen et al. (2024), attackers could exploit the open editing nature of these knowledge bases to inject malicious content that may alter the LLM’s outputs.
@@
This paper reveals a more realistic knowledge poisoning attack against RAG systems that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions involving complex relationships between multiple elements. As shown in Figure 1, such single document poisoning faces significant challenges. When the malicious attacker constructs a targeted document replacing “Kosovo” with “Nagorno-Karabakh”, the attack fails because: 1) The LLM prefers answers that appear across multiple retrieved documents rather than from a single source, as indicated by the multiple documents showing repeated mentions of “Kosovo”; 2) The LLM favors answers that align with its internal knowledge - in this case, its embedded knowledge correctly states that “Kosovo’s declaration did not violate international law”; 3) Simply injecting the question context into a single document creates unnatural content patterns that reduce the document’s credibility during retrieval and reasoning. These challenges make single document poisoning particularly difficult, as the attack must overcome both the LLM’s preference for consensus across multiple documents and alignment with its internal parameterized knowledge.
@@
To address these challenges, we propose AuthChain, a novel single document poisoning attack method that stealthily executes knowledge poisoning in RAG scenarios. Our approach follows three key principles for crafting an effective malicious document, with each progressively strengthens the attack. First, ensuring visibility. The poisoned document must stand out among vast external information sources. We achieve this through precise alignment with the question’s underlying objective. When external knowledge perfectly mirrors the core intent of a query, retrievers naturally prioritize it in their rankings, while LLMs tend to focus more on such intent-aligned information during their reasoning process. Second, guaranteeing retrievability and competitiveness. Even for complex queries involving multiple elements with complex relations, the document must remain a top candidate and outperform other retrieved knowledge. We accomplish this by structuring it as a self-contained evidence chain, preserving all key question elements and their logical relationships. This complete chain structure not only boosts retrieval rankings but also makes our content more compelling than fragmented knowledge pieces that only partially match the question’s logic. Third, overcoming LLMs’ internal knowledge bias. When LLMs consider multiple information sources, they tend to favor external knowledge that aligns with their internal knowledge. To counteract this bias, we strategically incorporate authority signals into our document, such as endorsements from authoritative institutions and recent timestamps. These authority signals help position our content as a more current and authoritative source compared to LLMs’ static internal knowledge, effectively overcoming their inherent preference for internally knowledge.
@@
2 Related Work
2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) has emerged as a crucial technique to enhance Large Language Models (LLMs) by addressing their hallucination issues when handling real-time and domain-specific questions Ye et al. (2023). The RAG architecture consists of three core components: a knowledge database, a retriever, and an LLM Zou et al. (2024). The knowledge database contains text chunks from various sources like Wikipedia, news articles, and financial documents Craswell et al. (2020); Loukas et al. (2023).
The RAG process operates in two main steps. First, during knowledge retrieval, the system employs two encoders (question encoder and text encoder) to convert both the query and documents into embedding vectors. The retriever then identifies the most relevant documents using similarity metrics . Second, in the answer generation phase, the LLM incorporates these retrieved documents as context to generate responses.
2.2 Existing Attacks on RAG
@@ 3 Methodology
To achieve effective single document poisoning in RAG, we propose AuthChain. Unlike previous approaches that require multiple poisoned documents or only effectively work for simple queries, AuthChain aims to create a malicious document that can not only be successfully retrieved among vast external knowledge sources, but also effectively compete with both authentic documents and LLMs’ internal knowledge, even for complex multi-hop questions. Our key insight is that the document’s influence can be progressively enhanced through three aspects: maximizing visibility during retrieval by precisely aligning with the question’s intent, enhancing both retrievability and competitiveness through complete evidence chains that maintain high retrieval rankings while outperforming fragmented external knowledge, and overcoming LLMs’ internal knowledge through authority reinforcement.
As shown in Figure 2, we implement these design principles through three main stages: (1) Intent-Based Content Generation focuses on maximizing document visibility by extracting three key features from the input question (the intent, key elements, and their relationships) and generates intent-based content. (2) CoE Content Generation aims to maintain high retrievability by constructing complete evidence chains. Using the extracted features and intent-based content, it generates Chain-of-Evidence (CoE) content that fully preserves the question’s semantic structure by covering its core objective, all key elements, and their relationships. (3) Authority Content Generation enhances document trustworthiness by incorporating domain-specific authority signals. Building upon the intent-based content, it creates authoritative content by incorporating institutional affiliations and recent timestamps, while maintaining professional formatting consistent with authoritative sources.
@@ 3.1 Intent-Based Content Generation
The example of question-derived features are presented in Appendix A. To effectively extract these features, AuthChain employs an LLM-based extraction approach enhanced with few-shot learning. Building upon the prompt template from  Li et al. (2023), we incorporate 5 carefully selected examples to improve extraction performance. The detailed prompt template and examples are provided in Appendix D.
Given the extracted intent, targeted question and answer, AuthChain prompts an intent agent to generate intent-based content. Such intent-driven content naturally aligns with the question’s essential requirements, which helps it achieve higher retrieval rankings and receive increased attention during LLM reasoning processes. Besides, it guides the subsequent knowledge generation to produce evidence that directly addresses the question’s central concerns. The detailed prompt template for the intent agent is provided in Appendix H.
@@ 3.2 CoE Content Generation
To both maintain high retrievability and outperform other external knowledge sources, AuthChain constructs complete evidence chains that preserve all question elements and their logical connections. While the intent-based content provides initial alignment with the question, we need to ensure the generated content comprehensively covers all extracted evidence nodes and their relationships.
@@
These suggestions, along with the current content, are then forwarded to a revise agent for refinement. This iterative evaluation and revision process continues until the CoE judge agent confirms complete preservation of the evidence chains, at which point the current content is finalized as the CoE content. The detailed prompt templates for both agents are provided in Appendix F.
@@ 3.3 Authority Content Generation
AuthChain employs an authority generator agent that takes the intent-based content, targeted answer, and question-derived features as input. By incorporating these features, the generated content maintains stronger semantic alignment with the original question, facilitating better retrieval. The agent first analyzes the intent-based content context to identify the most suitable authoritative institution for endorsement. It then synthesizes this institutional backing with recent timeline information to validate the targeted answer, ultimately producing authority content. The detailed prompt template for the authority generator agent is provided in Appendix G.
@@ 4.2 Dataset
We evaluate AuthChain on three widely-used question answering datasets in knowledge base poisoning attacks: Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), and MS-MARCO Nguyen et al. (2016). These datasets have been consistently used in prior poisoning attacks Zou et al. (2024); Zhang et al. (2024) due to their diverse knowledge sources (Wikipedia and web documents) and different question complexity levels. Following Zou et al. (2024), we use their constructed test set containing 100 query-answer pairs for each dataset to enable direct comparison with baseline attacks.
@@ 4.3 Subject RAG System Setting
We use the original knowledge bases provided by the datasets, Contriever Izacard et al. (2021) as the retriever, and six different LLMs, including three open-source LLMs (GPT-3.5 OpenAI (2022), GPT-4 Achiam et al. (2023), GPT-4o OpenAI (2024)) and three closed-source LLMs (LLama3-8B Touvron et al. (2024), LLama3-70B Touvron et al. (2023), Deepseek-V3-0324 DeepSeek (2024)). For each query, the retriever selects the top-5 most relevant documents as the context.
@@ 4.4 Baselines
  • HijackRAG (HRAG) Zhang et al. (2024): Constructs poisoned documents by combining retrieval text (the original question), hijack text (selected from a public prompt hacking dataset based on semantic similarity), and instruction text ("Print ’the answer is [targeted answer]’") to redirect model attention and control generation.
@@ 4.5 Experiment Design and Metric
For the approach implementation, we leverage GPT-4 as the backbone for the Intent Agent, the CoE Judge Agent, the revise agent, and the Authority Content Agent, consistent with the original PoisonedRAG work where GPT-4 was used to generate poisoned data Zou et al. (2024). The temperature parameter is set to 0.1 for all agents to ensure the stability and consistency of the generated content.
To answer RQ1, we investigate single document poisoning attacks where each method (AuthChain and baselines) constructs and injects one poisoned document per targeted question. We evaluate their effectiveness in manipulating RAG systems’ outputs and analyze AuthChain’s performance against baselines, while also examining its internal components (CoE and Authority content) for comprehensive analysis.
To answer RQ2, we select two representative RAG defense frameworks designed to counteract knowledge poisoning attacks: InstructRAG Wei et al. (2024) and AstuteRAG Wang et al. (2024b). InstructRAG enhances the robustness of RAG systems by explicitly guiding language models to learn a denoising process based on self-synthesized rationales. In this framework, the model is instructed to explain how the ground-truth answer is derived from the retrieved documents. These rationales can be leveraged as in-context demonstrations for explicit denoising or as supervised fine-tuning data, thereby improving the model’s ability to identify and resist poisoned or misleading knowledge in the retrieval set. AstuteRAG improves the robustness of RAG systems against imperfect or malicious retrieval by analyzing conflicts between the LLM’s internal knowledge and external sources. It adaptively extracts key information from internal knowledge, integrates it with retrieved content, and produces answers based on source reliability. This method has shown strong effectiveness in detecting and mitigating knowledge poisoning attacks. For fair comparison, we constrain all attacks (both AuthChain and baselines) to inject only a single poisoned document, and evaluate them under these two defense frameworks against the clean setting where no poisoned document is injected.
To answer RQ3, we conduct two experiments: (1) Authority Setting: Our goal is to examine whether authority-enhanced documents can influence LLM decisions even when they conflict with the LLM’s internal knowledge. To create an effective test environment, we need cases where LLMs have internal knowledge about the answers. We sample 600 QA pairs from HotpotQA and identify 118 questions that GPT-3.5 can correctly answer without external retrieval, indicating strong internal knowledge. These questions serve as ideal test cases since later GPT models also retain this internal knowledge. For these questions, we first create poisoned documents by modifying the correct answers in authentic documents to incorrect ones (Raw documents), then enhance these poisoned documents with authority signals using AuthChain. To simulate real-world scenarios where correct and incorrect information coexist, we gradually introduce authentic documents containing correct answers, creating mixed knowledge bases with correct document proportions (CDP) of 0.5, 0.67, and 0.75.
@@
We evaluate AuthChain using four metrics: Attack Success Rate (ASR), Retrieval Success Rate (RSR), Perplexity (PPL), and Accuracy (ACC). For evaluating the poisoning effectiveness (RQ1), we measure ASR as the proportion of questions where the LLM’s output contains the answer from the poisoned target document. Following previous works Rizqullah et al. (2023); Huang et al. (2023), we determine the presence of target answers using substring matching. We also examine RSR, which represents the proportion of poisoned target documents successfully retrieved among the top-5 documents, and PPL (calculated using GPT-2 Radford et al. (2019)) to measure text fluency where higher values indicate less natural text. For defense evaluation (RQ2), we compare both ASR and ACC, where ACC reflects the proportion of questions where the LLM’s response contains the correct answer. In RQ3, we investigate whether authority-enhanced documents can overcome LLMs’ internal knowledge bias by comparing ASR across different CDP, and whether CoE-structured documents are more influential than raw documents when competing with conflicting information by comparing ACC across different PDP.
@@ 5.1 Answering RQ1
We evaluate the effectiveness of AuthChain against six LLMs (GPT-3.5, GPT-4, GPT-4o, Llama3-8B, Llama3-70B, and DeepSeek-V3) across three widely-used datasets (HotpotQA, MS-MARCO, and NQ). Table 1 compares the Attack Success Rate (ASR) of AuthChain with both external baselines (PRAG and HRAG) and internal components (CoE and Authority content), along with key metrics including Retrieval Success Rate (RSR) and perplexity (PPL).
AuthChain achieves an average ASR of 87.0%, 81.5%, and 77.8% on HotpotQA, MS-MARCO, and NQ respectively, surpassing PRAG (21.7%-36.7%) and HRAG (31.0%-46.5%) by a significant margin. The results demonstrate the superior effectiveness of AuthChain on the single document poisoning scenario, which stems from our principle-guided content generation that better aligns with questions’ logical and semantic structure, further strengthened by authority expressions that enhance content credibility.
Specifically, while baselines resort to directly copying questions into poisoned documents, AuthChain dynamically generates content by incorporating key question elements through a principled three-stage design. The intent-based content generation successfully aligns with the question objective, while CoE content generation strengthens the logical alignment with question through constructed complete evidence chains. The effectiveness is further enhanced by our authority content generation which effectively leverages LLMs’ trust in authoritative sources. This structured approach generates more natural and authentic content while maintaining competitive retrievability, achieving only an average 2.3% decrease in RSR compared to PRAG and 5.0% higher than HRAG. The superior content quality is quantitatively validated by significantly lower perplexity scores (average PPL of 33.4 versus 69.9 for PRAG and 381.1 for HRAG).
Our component analysis reveals the effectiveness of both CoE and Authority content. Through an efficient generation process where judge agent and revise agent iterate only 1.3 times on average, our CoE content constructs complete evidence chains that preserve question structure, achieving 10.2%-18.8% higher ASR than baselines. This demonstrates its ability to efficiently compete with fragmented external knowledge while maintaining strong retrievability, with RSR only 4.0% lower than PRAG but 4.4% higher than HRAG. The Authority content shows 10.0%-18.6% ASR improvement through incorporating domain-specific authority signals, revealing LLMs’ susceptibility to authority bias across different architectures and scales. When combined, AuthChain surpasses CoE and authority content by 18.4% and 18.6% respectively, indicating strong complementarity: CoE content ensures competitive retrieval and dominance over external knowledge, while Authority content enhances trustworthiness to overcome LLMs’ internal knowledge.
Besides, to alleviate the computational costs of using GPT-4 based agents, we explore open-source LLMs as alternative agents in Appendix E. The results show these alternatives achieve comparable performance while being more cost-effective.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and found widespread applications in daily life. However, despite their impressive abilities, LLMs still face challenges such as outdated knowledge, hallucination, adversarial attacks, and jailbreak vulnerabilities as knowledge continues to evolve Achiam et al. (2023); Touvron et al. (2023); Anil et al. (2023); Jia et al. (2024); Teng et al. (2024); Lu et al. (2025); Guo et al. (2025). To address these limitations, researchers have introduced the Retrieval-Augmented Generation (RAG) approach to improve LLMs Tu et al. (2024); Zhao et al. (2024). This technology has been widely adopted in both industrial applications and academic research Al Ghadban et al. (2023); Wang et al. (2024a); Loukas et al. (2023); Kumar et al. (2023); Prince et al. (2024).
@@
Although RAG significantly improves the performance of the LLM response, it introduces potential security vulnerabilities. The security of RAG systems is influenced by both the inherent safety mechanisms of LLMs Tan et al. (2024); Xue et al. (2024); Chaudhari et al. (2024); Huang et al. (2024); Yang et al. (2025) and the potential manipulation of external context Zou et al. (2024); Zhang et al. (2024). Attackers can attempt to mislead LLMs into generating incorrect outputs by poisoning knowledge bases with malicious content. This attack surface is particularly concerning as knowledge bases are often the most accessible component of RAG systems. For example, when RAG systems utilize public resources like Wikipedia Thakur et al. (2021) for current events information Chen et al. (2024), attackers could exploit the open editing nature of these knowledge bases to inject malicious content that may alter the LLM’s outputs.
@@
This paper reveals a more realistic knowledge poisoning attack against RAG systems that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions involving complex relationships between multiple elements. As shown in Figure 1, such single document poisoning faces significant challenges. When the malicious attacker constructs a targeted document replacing “Kosovo” with “Nagorno-Karabakh”, the attack fails because: 1) The LLM prefers answers that appear across multiple retrieved documents rather than from a single source, as indicated by the multiple documents showing repeated mentions of “Kosovo”; 2) The LLM favors answers that align with its internal knowledge - in this case, its internal knowledge correctly states that “the declaration of independence of Kosovo”; 3) Simply injecting the question into a single document creates unnatural content patterns that reduce the document’s credibility during retrieval and reasoning. These challenges make single document poisoning particularly difficult, as the attack must overcome both the LLM’s preference for consensus across multiple documents and alignment with its internal parameterized knowledge.
@@
To address these challenges, we propose AuthChain, a novel single document poisoning attack method that stealthily executes knowledge poisoning in RAG scenarios. Our approach follows three key principles for crafting an effective malicious document, with each progressively strengthens the attack. First, ensuring visibility. The poisoned document must stand out among vast external information sources. We achieve this through precise alignment with the question’s underlying objective. When external knowledge perfectly mirrors the core intent of a question, retrievers naturally prioritize it in their rankings, while LLMs tend to focus more on such intent-aligned information during their reasoning process. Second, guaranteeing retrievability and competitiveness. Even for complex queries involving multiple elements with complex relations, the document must remain a top candidate and outperform other retrieved knowledge. We accomplish this by structuring it as a self-contained evidence chain, preserving all key question elements and their logical relationships. This evidence chain structure not only boosts retrieval rankings but also makes our content more compelling than fragmented knowledge pieces that only partially match the question’s logic. Third, overcoming LLMs’ internal knowledge bias. When LLMs consider multiple information sources, they tend to favor external knowledge that aligns with their internal knowledge. To counteract this bias, we strategically incorporate authority signals into our document, such as endorsements from authoritative institutions and recent timestamps. These authority signals help position our content as a more current and authoritative source compared to LLMs’ static internal knowledge, effectively overcoming their inherent preference for internally knowledge.
@@
2 Existing Attacks on RAG
@@ 3 Methodology
To achieve effective single document poisoning in RAG, we propose AuthChain. Our key insight is that the document’s influence can be progressively enhanced through three aspects: maximizing visibility during retrieval by precisely aligning with the question’s intent, enhancing both retrievability and competitiveness through self-contained evidence chains that maintain high retrieval rankings while outperforming fragmented external knowledge, and overcoming LLMs’ internal knowledge through authority reinforcement.
As shown in Figure 2, we implement these design principles through three main stages: (1) Intent-Based Content Generation focuses on maximizing document visibility by extracting three key features from the input question (the intent, key elements, and their relationships) and generates intent-based content. (2) CoE Content Generation aims to maintain high retrievability and competitive advantage over other knowledge sources by constructing self-contained evidence chains. Using the extracted features and intent-based content, it generates Chain-of-Evidence (CoE) content that fully preserves the question’s semantic structure by covering its core objective, all key elements, and their relationships. (3) Authority Content Generation enhances document trustworthiness by incorporating domain-specific authority signals. Building upon the intent-based content, it creates authoritative content by incorporating institutional affiliations and recent timestamps, while maintaining professional formatting consistent with authoritative sources.
@@ 3.1 Intent-Based Content Generation
The example of question-derived features are presented in Appendix A. To effectively extract these features, AuthChain employs an LLM-based extraction approach enhanced with few-shot learning. Building upon the prompt template from  Li et al. (2023), we incorporate 5 carefully selected examples to improve extraction performance. The detailed prompt template and examples are provided in Appendix C.
Given the extracted intent, targeted question and answer, AuthChain prompts an intent agent to generate intent-based content. The agent is instructed to generate content that not only provides the target answer, but also explicitly incorporates the question’s intent in the generated text. By formulating prompts that emphasize both answer generation and intent integration, the agent produces content that naturally aligns with the question’s essential objective, which helps it achieve higher retrieval rankings and receive increased attention during LLM reasoning processes. The detailed prompt template for the intent agent is provided in Appendix F.
@@ 3.2 CoE Content Generation
To both maintain high retrievability and outperform other external knowledge sources, AuthChain constructs self-contained evidence chains that preserve all question elements and their logical connections. While the intent-based content provides initial alignment with the question, we need to ensure the generated content comprehensively covers all extracted evidence nodes and their relationships.
@@
These suggestions, along with the current content, are then forwarded to a revise agent for refinement. This iterative evaluation and revision process continues until the CoE judge agent confirms complete preservation of the evidence chains, at which point the current content is finalized as the CoE content. The detailed prompt templates for both agents are provided in Appendix G.
@@ 3.3 Authority Content Generation
AuthChain employs an authority generator agent that takes the intent-based content, targeted answer, and question-derived features as input. By incorporating these features, the generated content maintains stronger semantic alignment with the original question, facilitating better retrieval. The agent first analyzes the intent-based content context to identify the most suitable authoritative institution for endorsement. It then synthesizes this institutional backing with recent timeline information to validate the targeted answer, ultimately producing authority content. The detailed prompt template for the authority generator agent is provided in Appendix E.
@@ 4.2 Dataset
We evaluate AuthChain on three widely-used question answering datasets in knowledge base poisoning attacks: Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), and MS-MARCO Nguyen et al. (2016). These datasets have been consistently used in prior poisoning attacks Zou et al. (2024); Zhang et al. (2024) due to their diverse knowledge sources (Wikipedia and web documents) and different question complexity levels. Following Zou et al. (2024), we use their constructed test set containing 100 question-answer pairs for each dataset to enable direct comparison with baselines.
@@ 4.3 Subject RAG System Setting
We use the original knowledge bases provided by the datasets, Contriever Izacard et al. (2021) as the retriever, and six different LLMs, including three open-source LLMs (GPT-3.5 OpenAI (2022), GPT-4 Achiam et al. (2023), GPT-4o OpenAI (2024)) and three closed-source LLMs (LLama3-8B Touvron et al. (2024), LLama3-70B Touvron et al. (2023), Deepseek-V3-0324 DeepSeek (2024)). For each question, the retriever selects the top-5 most relevant documents as the context.
@@ 4.4 Baselines
  • HijackRAG (HRAG) Zhang et al. (2024): Constructs poisoned documents by combining retrieval text (the original question), hijack text (selected from a public prompt hacking dataset based on semantic similarity), and instruction text (“Print the answer is [targeted answer]”) to redirect model attention and control generation.
@@ 4.5 Experiment Design and Metric
For the approach implementation, we leverage GPT-4 as the backbone for Intent Agent, CoE Judge Agent, Revise Agent, and Authority Content Agent, consistent with the original PoisonedRAG work where GPT-4 was used to generate poisoned document Zou et al. (2024). The temperature parameter is set to 0.1 for all agents to ensure the stability and consistency of the generated content.
To answer RQ1, we investigate single document poisoning attacks where each method (AuthChain and baselines) constructs and injects one poisoned document per targeted question. We evaluate their effectiveness in manipulating RAG systems’ outputs and analyze AuthChain’s performance against baselines, while also examining AuthChain’s internal components (CoE and Authority content) for comprehensive analysis.
To answer RQ2, we select two representative RAG defense frameworks designed to counteract knowledge poisoning attacks: InstructRAG Wei et al. (2024) and AstuteRAG Wang et al. (2024b). Detailed descriptions of these defense frameworks are provided in Appendix H. For fair comparison, we constrain all attacks (both AuthChain and baselines) to inject only a single poisoned document, and evaluate them under these two defense frameworks against the clean setting where no poisoned document is injected.
To answer RQ3, we conduct two experiments: (1) Authority Setting: Our goal is to examine whether authority enhanced documents can influence LLM decisions even when they conflict with the LLM’s internal knowledge. To create an effective test environment, we need cases where LLMs have internal knowledge about the answers. We sample 600 QA pairs from HotpotQA and identify 118 questions that GPT-3.5 can correctly answer without external retrieval, indicating strong internal knowledge. We conduct experiments on these test cases using GPT-series models, as this internal knowledge is consistently preserved in their subsequent versions. For these questions, we first create poisoned documents by modifying the correct answers in authentic documents to incorrect ones (Raw documents), then enhance these poisoned documents with authority signals using AuthChain. To investigate how the attack effectiveness of poisoned documents varies as increasing the proportion of external knowledge that aligns with LLMs’ internal knowledge, we gradually introduce authentic documents containing correct answers, creating mixed knowledge bases with correct document proportions (CDP) of 0.5, 0.67, and 0.75.
@@
We evaluate AuthChain using four metrics: Attack Success Rate (ASR), Retrieval Success Rate (RSR), Perplexity (PPL), and Accuracy (ACC). For evaluating the poisoning effectiveness (RQ1), we measure ASR as the proportion of questions where the LLM’s output contains the answer from the poisoned target document. Following previous works Rizqullah et al. (2023); Huang et al. (2023), we determine the presence of target answers using substring matching. We also examine RSR, which represents the proportion of poisoned target documents successfully retrieved among the top-5 documents, and PPL (calculated using GPT-2 Radford et al. (2019)) to measure text fluency where higher values indicate less natural text. For defense evaluation (RQ2), we compare both ASR and ACC, where ACC reflects the proportion of questions where the LLM’s response contains the correct answer. In RQ3, we investigate whether authority enhanced documents can overcome LLMs’ internal knowledge bias by comparing ASR across different CDP, and whether CoE-structured documents are more influential than raw documents when competing with conflicting information by comparing ACC across different PDP.
@@ 5 Result
5.1 Answering RQ1
We evaluate the effectiveness of AuthChain against six LLMs (GPT-3.5, GPT-4, GPT-4o, Llama3-8B, Llama3-70B, and DeepSeek-V3-0324) across three widely-used datasets (HotpotQA, MS-MARCO, and NQ). Table 1 compares the Attack Success Rate (ASR) of AuthChain with both baselines (PRAG and HRAG) and internal components (CoE and Authority content), along with key metrics including Retrieval Success Rate (RSR) and perplexity (PPL).
AuthChain achieves an average ASR of 87.0%, 81.5%, and 77.8% on HotpotQA, MS-MARCO, and NQ respectively, surpassing PRAG (21.7%-36.7%) and HRAG (31.0%-46.5%) by a significant margin. The results demonstrate the superior effectiveness of AuthChain on the single document poisoning scenario, which stems from our CoE content generation that better aligns with questions’ logical and semantic structure, further strengthened by authority expressions that enhance content credibility.
Besides, while baselines resort to directly copying questions into poisoned documents, AuthChain dynamically generates content by incorporating self-contained evidence chains and authoritative signals that synthesize contextually-appropriate institutional endorsements with temporal validations to support the targeted answer. Therefore, AuthChain generates more natural and authentic content while maintaining competitive retrievability, achieving only an average 2.3% decrease in RSR compared to PRAG and 5.0% higher than HRAG. The superior content quality is quantitatively validated by significantly lower perplexity scores (average PPL of 33.4 versus 69.9 for PRAG and 381.1 for HRAG).
Our component analysis reveals the effectiveness of both CoE and Authority content. Through an efficient generation process where judge agent and revise agent iterate only 1.3 times on average, the CoE content constructs self-contained evidence chains that preserve question structure, achieving 10.2%-18.8% higher ASR than baselines. This demonstrates its ability to efficiently compete with fragmented external knowledge while maintaining strong retrievability, with RSR only 4.0% lower than PRAG but 4.4% higher than HRAG. The Authority content shows 10.0%-18.6% ASR improvement through incorporating domain-specific authority signals, revealing LLMs’ susceptibility to authority bias across different architectures and scales. When combined, AuthChain surpasses CoE and authority content by 18.4% and 18.6% respectively, indicating strong complementarity: CoE content ensures competitive retrieval and dominance over external knowledge, while Authority content enhances trustworthiness to overcome LLMs’ internal knowledge.
Additionally, to alleviate the computational costs of using GPT-4 based agents, we explore open-source LLMs as alternative agents in Appendix D. The results show these alternatives achieve comparable performance while being more cost-effective.